OpenStudy (anonymous):

3 -- (30x) 5 I got 150x ----- 3 Is that the final answer, or can i simplify it more?

OpenStudy (anonymous):

Is this what you mean: \[\frac{ 3 }{ 5 } (30x) \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 } \times (30x)\] MULTIPLY 3 BY 30X IN THE NUMERATOR \[\frac{ 3 \times 30x }{ 5 }\] \[\frac{ 90x }{ 5 }\] REDUCE THE EXPRESSION BY REMOVING A FACTOR OF 5 FROM THE NUMERATOR AND DENOMINATOR \[18x \]

OpenStudy (anonymous):

Oh..Thank you! I cross multiplied b4

OpenStudy (anonymous):

Well you can still cross multiply, but like this: \[\frac{ 3 }{ 5 } \times (30x) \] \[\frac{ 3 }{ 5 } \times \frac{ 30x }{ 1 }\] \[\frac{ 3 \times 30x }{ 5 \times 1 }\] \[\frac{ 90x }{ 5 }\] \[18x \]

OpenStudy (anonymous):

Wait, never mind; cross multiplying will not work.